OpenStudy (anonymous):

Stanley and Phyllis leave the office and travel in opposite directions. Stanley drives 6 MPH slower than Phyllis and after 1 hr they are 76 miles apart. How fast was each person driving?

OpenStudy (wolf1728):

S = P - 6 (both rates are in mph) Together they both travel 76 miles in one hour. S + P = 76 P -6 + P = 76 2 P = 82 So Phyllis drives at 41 mph and Stanley travels at 35 mph

OpenStudy (anonymous):

Thank you ! @wolf1728